OpenStudy (anonymous):

A girl is now one-fourth as old as her father, and in seven years, she will be one-half as old as her father was twelve years ago. What are her and her father's present ages? father's age = 20; daughter's age = 5 father's age = 52; daughter's age = 13 father's age = 76; daughter's age = 19

OpenStudy (anonymous):

so, if f is the father's age, and d the daughter's, d=f/4 and d+7=(f-12)/2 then by substituting f/4 into the second equation, f/4+7=f/2-6 f/4=13 f=52 and we're done!